# Liwen Wang1,2,  Yuanyuan Yuan2,  Ao Sun2,  Zongjie Li2,  Pingchuan Ma2,  Daoyuan Wu2,  Shuai Wang2

### v2

@@ 6 Human Studies
We form a group of 20 participants to evaluate the ten mutations in Table 2 according to the alignment, faithfulness, and validity requirements (discussed in Sec. 5.1). The effectiveness of these mutations in triggering VDL faults is evaluated using quantitative evaluations in Sec. 7.3. Since evaluating if a mutation fulfills these requirements needs inspecting how the changed semantics (both anticipated and unanticipated) may affect the downstream VDL testing and analysis, we hire 20 Ph.D. students experienced in software testing and VDL system in our human study.
6.1 Human Study Design
For each mutation listed in Table 2, we prepare 50 seed images to mutate. That is, for the six traditional mutations, each of them has $50\times 5=250$ mutated images, and for the four MLLM-enabled mutations, each of them has $50\times 4=200$ mutated images.
@@ 6.1 Human Study Design
The common practice, as in most existing MLLM works [68, 53], is letting humans rank different mutated images. However, we observed that MLLMs can fail in implementing a mutation or generating valid images, which cannot be reflected from the ranking results. To address this, we design the questions as follows. Each question in our human study is formed by a seed image and one of its mutated variants (using different implementations), and a participant is asked to give a score (0 to 5) for the alignment, faithfulness, and validity of a mutated image. A zero score indicates that the mutated image fails to satisfy one requirement. For example, if the rotation mutation implemented using a MLLM does not rotate an image, the alignment score should be zero. We use scores 1-5 to assess each requirement because one mutation has maximal five different implementations (i.e., one traditional method + four MLLMs); the five different choices can distinguish different implementations.
@@
We let each participant evaluate one of the ten mutations, in the sense that interchangeably evaluating different mutations may confuse the participants, especially for subtle mutations that require careful comparison. Therefore, each participant answers 200 (for MLLM-enabled mutations) or 250 (for traditional mutations) questions. To reduce personal preferences of participants, each mutation is assigned to two different participants. The orders of the questions are shuffled to avoid potential bias due to question orders.
@@
Before starting the human study, we prepare a 30-minute teaching for each participant. The teaching primarily aims to introduce the three requirements (i.e., alignment, faithfulness, and validity) with different successful and failed examples. It also gives participants suggestions on giving scores. For instance, zero indicates an unsatisfied requirement, and the scores should be distinguishable. Our human study is conducted using the Amazon Mechanical Turk platform [1]. We do not set time limit for each question because we expect participants to carefully assess the mutated images. The participant can pause and resume the evaluation at any time to avoid fatigue. The average time to finish all evaluation questions (i.e., 200 or 250) is around two hours as reported by participants.
@@
6.2 Human Study Results
Table 3 reports the average score over all 200 $\times$ 2 or 250 $\times$ 2 samples for each mutation. Since a zero score indicates that a mutated image fails to satisfy one requirement, we deem scores within $[0,1)$ as failing and mark them in red. We view scores larger than two as satisfying the requirement. Specifically, scores within $(2,4]$ are marked in blue, whereas scores within $(4,5]$ are marked in green to highlight the best method for one mutation.
@@ 6.2 Human Study Results
Contrast and Rotation: Our results show that MLLMs are unable to implement pixel- and geometrical-level mutations, as supported by the red scores under the alignment evaluation. Note that pixel- and geometrical-level mutations directly operate on image pixels, whereas the remaining mutations focus on image features (i.e., image representations obtained using neural networks). MLLMs also primarily accept image features, leaving lower-level image semantics untouched. This explains why MLLMs fail in implementing pixel- and geometrical-level mutations, as images before and after these mutations often do not lead to notable changes in image features. Recent MLLM papers have similar observations: MLLMs perform worse in the spatial reasoning of image contents [49, 7, 12].
@@
Artifying: MLLMs outperform traditional methods in transferring the artistic style of images. Past methods often extract the artistic style from one or a few target images, whose understanding of the target style may be limited. Moreover, traditional style transfer may fail if contents in the seed image and the target have distinct geometrical structures. These two limitations explain why the alignment score of traditional method is unsatisfactory (i.e., only 1.7). MLLMs encode general knowledge and can better comprehend the target style from their immense amounts of training data.
@@
SnowG and SnowD: InstP2P fails to transfer weather conditions for general images and driving scenes. We attribute the reason to the lack of fine-tuning data. To support transferring weather conditions, the fine-tuning in InstP2P requires mutation variants of the seed image that contain different weather conditions, and such data tuples are hard to prepare in large scale. Thus, the fine-tuned Diffusion model in InstP2P is unable to capture the weather conditions, transmitting the input image to the output in most cases, as indicated by the high faithfulness and validity score. The traditional method leverages the cycle-consistency to achieve weather transfer without requiring paired images as the guidance. This paradigm successfully transfers the weather conditions for diving scenes, resulting in a high alignment score. However, it fails to retain the faithfulness (i.e., a 0.2 score), such that mutated images do not preserve irrelevant semantics, as presented in Fig. 3. In addition, the cycle-consistency relies on the similar geometrical structure between the seed and target images; this explains why traditional method is inapplicable to mutate weather conditions in general images whose geometrical structures are largely distinct from driving scenes.
@@
Eyes: Traditional methods outperform MLLMs when implementing perceptual-level mutations. All MLLMs are unable to deliver satisfactory alignment scores. As in Fig. 3, all MLLMs do not close the eyes but replace them with new eyes in most cases. That is, these MLLMs understand that the mutation modifying eyes but cannot precisely mutate eyes. Moreover, recall that as mentioned in Sec. 3, the traditional implementation of Eyes requires specifying the localization of eyes. To make the comparison fair, we also provide the eye locations to MLLMs. However, as shown in Fig. 3, MLLMs fail to generate smoothly mutated images and the edited eyes look unnatural, leading to relatively lower validity scores.
@@
MLLM-Enabled (⑦-⑩): Our results show that only InstP2P can implement Background to a satisfactory level, whereas the other three MLLMs tend to not edit the seed image in Background, as indicated by the relatively high faithfulness and validity score (see examples in Fig. 3). However, the trend is on the opposite when implementing Clothes and Legs, where InstP2P has lower alignment scores but much higher faithfulness and validity scores. For Glasses, all MLLMs have satisfactory alignment, faithfulness, and validity scores, and InstP2P outperforms the other three MLLMs.
To understand these distinct results, we need to dig into the key difference between mutations ⑦-⑩. Holistically, Glasses and Background are more “silent” whereas Clothes and Legs are more “active”. Background focuses on the background, whereas the other three mutations edit the subjects in the seed image. Glasses can be viewed as the simplest form of semantic-additive mutation as it is similar to “inserting” a new object into the seed; it only requires locating the eyes and putting the glasses on without altering other content. Clothes and Legs, in contrast, require additionally modifying the remaining semantics for coherence consideration.
7 Quantitative Evaluation
We also conduct a large-scale quantitative evaluation to assess these input mutations using popular numerical metrics. While the human study results are sufficient to reflect the quality of MLLM-produced test inputs, we wish the quantitative evaluations, which are automated and larger-scale, can further justify the human study results and provide insights from different angles.
7.1 Alignment and Faithfulness Evaluation
Our evaluated mutations focus on different semantics and operate at distinct levels. Thus, the alignment and faithfulness evaluations rely on the implementation of a mutation. We separately design alignment and faithfulness evaluations for different mutations.
Table 4: Evaluation results of Contrast & Rotation.
Contrast: MSE Rotation: SSIM
LLaVA 3039.72 0.177
CogVLM 3112.78 0.178
GPT-4V 3034.89 0.175
InstP2P 2532.89 0.134
Contrast (①) and Rotation (②): Given that pixel- and geometrical-level mutations are explicitly defined as mathematical transformations, mutations implemented using traditional methods are the ground truth. Also, considering that these two types of mutations globally mutate all pixels, it is infeasible to separately evaluate the alignment and faithfulness using numerical metrics. Therefore, we jointly evaluate the alignment and faithfulness of Contrast and Rotation implemented using MLLMs by comparing their mutated images with those of traditional methods. Specifically, because Contrast uniformly edits all pixel values, we compute the Mean Squared Error (MSE) between pixel values of images produced by MLLMs and traditional methods. A lower MSE value indicates that the MLLM-produced images are closer to the ground truth. For Rotation, since it moves the locations of pixels, we use SSIM, a metric of image structure-wise similarity, to quantify the differences between images mutated by MLLMs and traditional methods. The SSIM value ranges within $[0,1]$ and a higher value indicates that the MLLM’s mutated image is more similar to the ground truth.
Results are in Table 4. For Contrast, the differences between MLLM’s mutation and the ground truth are high. As a reference, the average MSE between seed images and their mutated outputs using Contrast is 1167.16, showing that MLLM’s mutated images largely deviate from the expected mutation. The MSE of InstP2P is lower than other three MLLMs, which is aligned to our human study results, where InstP2P’s alignment score is 1.2 and other MLLMs’ alignment scores are $<$ 1. Rotation implemented using MLLMs are largely dissimilar to the ground truth (higher SSIM values denoting better results). The SSIM value of InstP2P is much lower than other MLLMs. This is aligned to our human study results: InstP2P generates invalid images under Rotation and results in much lower faithfulness and validity scores. Hence, we conclude that MLLMs can hardly support pixel- and geometrical-level mutations.
Table 5: Quantitative results of style-level mutations.
Artifying SnowG SnowD
Style loss* (Alignment) LLaVA 2.67 1.94 1.06
CogVLM 2.73 1.93 1.08
GPT-4V 2.62 2.12 1.13
InstP2P 2.60 1.70 1.03
Traditional 2.39 1.23 1.06
Content loss (Faithfulness) LLaVA 8.96 8.21 4.18
CogVLM 8.90 8.01 3.90
GPT-4V 8.25 7.98 4.12
InstP2P 4.92 2.43 1.03
Traditional 3.02 2.97 3.74
  • * The unit of style loss is 1e-4.
Style-level Mutations (③-⑤). Recall that as mentioned in Sec. 3.2, style-level mutations separate an image as the content and style and aim to edit the image’s style. Therefore, following existing works on style-level mutations, we evaluate the alignment and faithfulness of style-level mutations with the style loss $L_{\text{style}}$ and content loss $L_{\text{content}}$ [23], respectively. The $L_{\text{style}}$ prepares a set of images having the target “style” (e.g., Van Gogh’s paintings, or photos taken on snowy days) and extracts their style features using a pre-trained feature extractor. It then extracts the style feature from the mutated image and compares it with the target style features. A lower $L_{\text{style}}$ value indicates a better alignment. Similarly, the $L_{\text{content}}$ extracts and compares the content features of the seed image and the mutated image; a lower $L_{\text{content}}$ value indicates a better faithfulness.
As shown in Table 5, although the style loss is widely used in existing style-level mutation works for benchmark or optimization, it cannot notably distinguish results of different implementations and is unable to reveal our findings in the human studies. The content loss, in contrast, varies among different implementations of the same style-level mutation. For Artifying and SnowG, the content loss values of InstP2P and traditional method are largely lower than others. Similarly, SnowD implemented using InstP2P also has a much lower content loss, indicating the content of the seed image is mostly preserved in the mutated image. These results are consistent to the faithfulness evaluations in our human studies. Overall, InstP2P tends to not mutate the seed image when implementing all style-level mutations. Traditional methods hardly mutate the weather conditions of general images (SnowG) and only slightly apply artistic style to seed images (Artifying).
Table 6: Anticipated results of MLLM-enabled mutations.
@@ 7.1 Alignment and Faithfulness Evaluation
MLLM-enabled Mutations (⑦-⑩): Due to the lack of ground truth and reference images for mutations ⑦-⑩, we evaluate their alignment via CLIP [48] following recent MLLM papers [53, 68, 28]. The CLIP can obtain unified embeddings for image semantics and text descriptions, enabling their similarity measurement. We first prepare the textual description of the anticipated mutation result for each mutation, as shown in Table 6. For example, when mutating a dog image with Glasses, the mutated image should have “a dog with glasses”. We then use the CLIP to extract the embedding vectors for the mutated image and its corresponding textual description in Table 6; their cosine similarity is adopted to quantify the alignment of the mutation.
The faithfulness evaluation requires separating the local and fine-grained semantics such as eyes and legs from the image, which is challenging especially when the locations and postures of the dog are diverse. Therefore, we design a coarse-grained metric that only considers an image’s background and subject. Specifically, following [31], we employ the SoTA subject extractor InSPyReNet [38] to separate the subjects and background for the seed and mutated images. For Background, we use the similarity of the two subjects in the seed and mutated image as the faithfulness result, whereas for ⑧-⑩, we instead use the similarity of two backgrounds. Since semantic-additive mutations may alter the subject’s outline (e.g., adding clothes may change the dog’s shape), we compute the similarity as the cosine similarity between CLIP image embeddings of the extracted subjects/backgrounds.
Table 7: Alignment (left) and faithfulness (right) results of MLLM-enabled mutations.

### v3

@@ 6 Human Studies Setup
We form a group of 20 participants to evaluate the ten mutations in Table II according to the alignment, faithfulness, and validity requirements (discussed in Sec. 5.1). Since evaluating if a mutation fulfills these requirements needs inspecting how the changed semantics (both anticipated and unanticipated) may affect the downstream VDL testing and analysis, we hire 20 Ph.D. students experienced in software testing and VDL systems in our human study. For each mutation in Table II, we prepare 50 seed images to mutate. That is, for the six traditional mutations, each of them has $50\times 5=250$ mutated images, and for the four MLLM-enabled mutations, each of them has $50\times 4=200$ mutated images.
@@
The common practice, as in most existing MLLM works [63, 64], is letting humans rank different mutated images. However, we observed that MLLMs can fail in implementing a mutation or generating valid images, which cannot be reflected in the ranking results. To address this, we design the questions as follows. Each question in our human study is formed by a seed image and one of its mutated variants (using different implementations), and a participant is asked to give a score (0 to 5) for the alignment, faithfulness, and validity of a mutated image. A zero score indicates that the mutated image fails to satisfy one requirement. For example, if the rotation mutation implemented using a MLLM does not rotate an image, the alignment score should be zero. We use scores 1-5 to assess each requirement because one mutation has maximal five different implementations (i.e., one traditional method + four MLLMs); the five different choices should be distinguishable for different implementations.
@@
We let each participant evaluate one of the ten mutations, as interchangeably evaluating different mutations may confuse the participants, especially for subtle mutations that require careful comparison. Thus, each participant answers 200 (for MLLM-enabled mutations) or 250 (for traditional mutations) questions. To reduce personal preferences, each mutation is assigned to two different participants. The orders of the questions are shuffled to avoid potential bias due to question orders.
To rule out invalid evaluations (e.g., a participant misunderstands the requirement), following previous work [65], we insert sanity-check questions into each evaluation without notifying the participants. Specifically, we insert 10 questions where the “mutated” image is identical to the seed image, and the participant is expected to assign 0, 5, 5 scores for the alignment, faithfulness, and validity requirements, respectively. If a participant fails to pass 80% of the sanity-check questions, we will discard all his/her answered questions. It’s worth noting that all our participants pass the sanity check.
@@
Before starting the human study, we prepare a 30-minute teaching for each participant. The teaching primarily aims to introduce the three requirements (i.e., alignment, faithfulness, and validity) with different successful and failed examples. It also gives participants suggestions on giving scores. For instance, zero indicates an unsatisfied requirement, and the scores should be distinguishable. Our human study is conducted using the Amazon Mechanical Turk platform [30]. We do not set a time limit for each question because we expect participants to carefully assess the mutated images. The participant can pause and resume the evaluation at any time to avoid fatigue. The average time to finish all evaluation questions (i.e., 200 or 250) is around two hours as reported by participants.
7 Results and Findings.
This section presents the results of our human study and answers three RQs mentioned in Sec. 5.
@@
7.1 RQ1: Unifying and Generalizing Mutations
Table III reports the average score over all 200 $\times$ 2 or 250 $\times$ 2 samples for each mutation. Since a zero score indicates that a mutated image fails to satisfy one requirement, we deem scores within $[0,1)$ as failing and mark them in red. We view scores larger than two as satisfying the requirement. Specifically, scores within $(2,4]$ are marked in blue, whereas scores within $(4,5]$ are marked in green to highlight the best method for one mutation.
@@ 7.1 RQ1: Unifying and Generalizing Mutations
Contrast and Rotation. Our results show that MLLMs are unable to implement pixel- and geometrical-level mutations, as supported by the red scores under the alignment evaluation. Note that pixel- and geometrical-level mutations directly operate on image pixels, whereas the remaining mutations focus on image features (i.e., image representations obtained using neural networks). MLLMs also primarily accept image features, leaving lower-level image semantics untouched. This explains why MLLMs fail in implementing pixel- and geometrical-level mutations, as images before and after these mutations often do not lead to notable changes in image features. Recent MLLM papers have similar observations: MLLMs perform worse in the spatial reasoning of image contents [66, 67, 68].
@@
Artifying. MLLMs outperform traditional methods in transferring the artistic style of images. Past methods often extract the artistic style from one or a few target images, whose understanding of the target style may be limited. Moreover, traditional style transfer may fail if contents in the seed image and the target have distinct geometrical structures. These two limitations explain why the alignment score of the traditional method is unsatisfactory (i.e., only 1.7). MLLMs encode general knowledge and can better comprehend the target style from their immense amounts of training data.
@@
SnowG and SnowD. InstP2P fails to transfer weather conditions for general images and driving scenes. We attribute the reason to the lack of fine-tuning data. To support transferring weather conditions, the fine-tuning in InstP2P requires mutation variants of the seed image that contain different weather conditions, and such data tuples are hard to prepare on a large scale. Thus, the fine-tuned Diffusion model in InstP2P is unable to capture the weather conditions, transmitting the input image to the output in most cases, as indicated by the high faithfulness and validity score. The traditional method leverages the cycle-consistency to achieve weather transfer without requiring paired images as the guidance. This paradigm successfully transfers the weather conditions for diving scenes, resulting in a high alignment score. However, it fails to retain the faithfulness (i.e., a 0.2 score), such that mutated images do not preserve irrelevant semantics, as presented in Fig. 4. In addition, the cycle-consistency relies on the similar geometrical structure between the seed and target images; this explains why traditional method is inapplicable to mutate weather conditions in general images whose geometrical structures are largely distinct from driving scenes.
@@
Eyes. Traditional methods outperform MLLMs when implementing perceptual-level mutations. All MLLMs are unable to deliver satisfactory alignment scores. As in Fig. 4, all MLLMs do not close the eyes but replace them with new eyes in most cases. That is, these MLLMs understand that the mutation modifying eyes but cannot precisely mutate eyes. Moreover, recall that as mentioned in Sec. 4, the traditional implementation of Eyes requires specifying the localization of eyes. To make the comparison fair, we also provide the eye locations to MLLMs. However, as shown in Fig. 4, MLLMs fail to generate smoothly mutated images and the edited eyes look unnatural, leading to relatively lower validity scores.
Answer to RQ1: MLLMs cannot implement pixel- and geometrical-level mutations well, potentially due to the lost information of such low-level semantics when MLLMs extract image features. MLLMs are also unable to generalize perceptual-level mutations. However, MLLMs optimized with the bidirectional pipeline perform better in extending style-level mutations to more general scenarios. They also exhibit better resilience: since they do not modify the image generator, they will not generate invalid images in case they cannot achieve a mutation.
7.2 RQ2: Expanding Mutated Semantics
@@
MLLM-Enabled (⑦-⑩). As in Table III, we find that only InstP2P can implement Background to a satisfactory level, whereas the other three MLLMs tend not to edit the seed image in Background, as indicated by the relatively high faithfulness and validity score (see examples in Fig. 4). However, the trend is on the opposite when implementing Clothes and Legs, where InstP2P has lower alignment scores but much higher faithfulness and validity scores. For Glasses, all MLLMs have satisfactory alignment, faithfulness, and validity scores, and InstP2P outperforms the other three MLLMs.
Overall, these results show that MLLMs work well in semantic-replacement mutations, indicating MLLM’s high potential in expanding prior mutations with a new dimension. The varied results of different pipelines (see Sec. 5.2) reveal their specialties. Both optimized pipelines support subtle mutations to the subjects. However, the unidirectional pipeline is more suitable for background, while the bidirectional pipeline fits better when largely mutating subjects in images.
Testing Effectiveness. To further study whether MLLM-enabled mutations are useful in VDL testing, we compare their effectiveness in detecting VDL failures with traditional mutations. Contrast (①) and Rotation (②) are implemented with their original methods. Artifying (③) and SnowG (④) are implemented using MLLMs as MLLMs perform better than their original implementations. Note that SnowG denotes SnowD’s application in general images; SnowD is therefore omitted. We also omit Eyes in this evaluation since Eyes (⑥) is inapplicable to general images and MLLM also cannot generalize it according to results in Sec. 7.1.
VDL-based classification is selected as one representative VDL application. We choose three recent SoTA VDL-based classifiers, Vision Transformer (ViT) [69], SwinTransformer (SwinT) [70], and ConvNeXt [71]. These models are officially released by PyTorch and are trained with ImageNet.
[…]
Model Contrast Rotation Artifying SnowG Background Clothes Legs Glasses
ConvNeXt LLaVA N/A N/A 7.6% 4.8% 5.2% 7.7% 7.7% 2.8%
CogVLM 9.6% 4.8% 7.6% 8.6% 10.1% 4.5%
GPT4V 9.2% 3.6% 6.8% 5.7% 9.7% 1.6%
InstructP2P 6.4% 4.4% 11.2% 0.8% 1.2% 1.2%
Traditional 6.0% 0.4% 6.8% 14.4% N/A N/A N/A N/A
SwinTransformer LLaVA N/A N/A 8.0% 4.8% 5.2% 7.7% 7.3% 2.8%
CogVLM 9.6% 4.8% 7.2% 9.4% 10.9% 4.5%
GPT4V 9.2% 4.4% 7.2% 6.1% 9.3% 1.6%
InstructP2P 5.6% 4.0% 13.2% 2.0% 2.8% 1.6%
Traditional 6.8% 0.4% 8.4% 16.0% N/A N/A N/A N/A
VisionTransformer LLaVA N/A N/A 8.0% 4.8% 4.4% 6.1% 7.7% 3.2%
CogVLM 9.6% 5.2% 7.6% 9.8% 10.5% 4.9%
GPT4V 7.6% 4.4% 7.6% 5.3% 10.1% 0.8%
InstructP2P 7.2% 3.6% 12.0% 1.6% 2.0% 1.6%
Traditional 8.4% 0.4% 7.2% 16.0% N/A N/A N/A N/A
  • * N/A indicates that the mutation cannot be implemented using MLLMs.
[…]
Answer to RQ2: MLLMs are complementary to traditional mutations and expand their mutated semantics by uniquely enabling new semantic-replacement mutations. These mutations are also effective in detecting VDL faults. The unidirectional mutation pipeline of MLLMs performs better for background semantics, whereas the bidirectional pipeline is suitable for noticeable mutations on subject semantics.
7.3 RQ3: Validation Metrics
Aligned to the three aspects considered in our human study, we evaluate whether existing validation metrics can reflect the validity of test inputs generated by MLLM-enabled mutations and capture the varied extents of MLLM-enabled mutations.
7.3.1 Validity Metrics
Following existing works [72, 73], we take two popular metrics, Inception Score (IS) [49] and Frechet Inception Distance (FID) [50], to measure the validity of the mutated images. Both IS and FID use a set of real images as the reference, and compare the mutated images with these reference images to estimate how likely the mutated images are real. Lower FID values indicates better realism, whereas higher IS values indicates better realism.
TABLE V: Validity results of MLLM-enabled mutations.
LLaVA CogVLM GPT-4V InstP2P
IS 14.51 13.71 14.82 23.93
FID 88.09 89.65 88.67 93.95
As shown in Table V, IS values can roughly distinguish InstP2P from other MLLMs. While this is aligned with our human evaluation results, we expect finer-grained differences among different MLLMs can be accurately reflected. FID values, however, show contradictory results with IS values. Overall, observations in Table V indicate the inaccuracy of existing validity metrics when facing MLLM-enabled mutations. As pointed out by previous works [74, 75], FID and IS jointly consider the validity and diversity. This was not an issue previously because seed and mutated images are highly similar when applying most traditional mutations (e.g., Rotation). However, semantic-replacement mutations enabled by MLLMs bring non-determinism: multiple possible mutants can be generated when applying one mutation to a seed image. As a result, the validity metric’s results can be falsely improved by diversified mutants in MLLM-enabled mutations.
7.3.2 Embedding-Based Metrics
TABLE VI: Anticipated results of MLLM-enabled mutations.
@@ 7.3.2 Embedding-Based Metrics
As mentioned in Sec. 3.2, for mutations lacking explicit formulas (i.e., style-level, perceptual-level, and MLLM-enabled mutations), existing works propose to measure them via image embeddings. Following recent MLLM papers [64, 63, 76, 51], we obtain image embeddings via the SoTA embedding model CLIP [15]. The CLIP can obtain unified embeddings for image semantics and text descriptions, enabling their similarity measurement. We first prepare the textual description of the anticipated mutation result for each mutation, as shown in Table VI. For example, when mutating a dog image with Glasses, the mutated image should have “a dog with glasses”. We then use the CLIP to extract the embedding vectors for the mutated image and its corresponding textual description in Table VI; their cosine similarity is adopted to quantify the alignment of the mutation.
The faithfulness evaluation requires separating the local and fine-grained semantics such as eyes and legs from the image, which is challenging especially when the locations and postures of the dog are diverse. Therefore, we design a coarse-grained metric that only considers an image’s background and subject. Specifically, following [77], we employ the SoTA subject extractor InSPyReNet [78] to separate the subjects and background for the seed and mutated images. For Background, we use the similarity of the two subjects in the seed and mutated image as the faithfulness result, whereas for ⑧-⑩, we use the similarity of two backgrounds for faithfulness. Since semantic-replacement mutations may alter the subject’s outline (e.g., adding clothes may change the dog’s shape), we compute the similarity as the cosine similarity between CLIP image embeddings of the extracted subjects/backgrounds.
TABLE VII: Alignment (left) and faithfulness (right) results of MLLM-enabled mutations.